OpenStudy (anonymous):

could you help me with one more thing? @satellite73

OpenStudy (anonymous):

sure if i know it

OpenStudy (solomonzelman):

you know it.

OpenStudy (anonymous):

how can I derivate double Angle Formulas from the sum of Two Angles?

OpenStudy (anonymous):

by changing \(2x\) in to \(x+x\) and using it i.e put \[\sin(x+y)=\sin(x)\cos(y)+\cos(x)\sin(y)\] and make it \[\sin(x+x)=...\]

OpenStudy (anonymous):

in simple english, where you see a \(y\) replace it by an \(x\)

OpenStudy (anonymous):

oh ok so I'm allowed to just do that?

OpenStudy (anonymous):

i give you permission, don't tell

OpenStudy (anonymous):

lol yes you get it essentially in one step

OpenStudy (anonymous):

haha alright thanks again :)

OpenStudy (anonymous):

yw since you probably have the double angle formula written down you can check that your answer is correct you will get 3 different ones for the double angle formula for cosine that one turns out to be most useful